OpenStudy (anonymous):

f'x=1/(2x^1/2) and f(1)=2, find f(1/2)

OpenStudy (anonymous):

How do I go about solving this? Do I find the antiderivative?

OpenStudy (anonymous):

Is this f'(x)=1/(2(x)^1/2) or f'(x)=1/(2x)^1/2 ?

OpenStudy (apples):

Yes, then you find the constant of integration by setting your antiderivative equal to 2 and evaluating where x = 1. Once you know the constant, you can evaluate f(1/2) like normal.

OpenStudy (anonymous):

sorry, it's 1/ ( 2 (x)^1/2)

OpenStudy (anonymous):

I found that the antiderivate is x^1/2, is that what you got

OpenStudy (anonymous):

Your antiderivative is going to be x^(1/2)+c

OpenStudy (apples):

Right, \[\sqrt{x} + C\]

OpenStudy (apples):

\[f(x) = \int \frac{1}{2\sqrt{x}} dx = \sqrt{x} + C\]\[\sqrt{1}+C=2\] Thus C is 1 and \[f(\frac{1}{2}) = \sqrt{\frac{1}{2}} + 1\]

OpenStudy (anonymous):

cool thanks a lot

OpenStudy (anonymous):

anyone wanna help for lim x->infinity x*sin(1/x), is sin(1/x) 0?